# Exploring Training and Inference Scaling Laws in Generative Retrieval

### v1

@@ Abstract.
Generative retrieval has emerged as a novel paradigm that leverages large language models (LLMs) to autoregressively generate document identifiers. Although promising, the mechanisms that underpin its performance and scalability remain largely unclear. We conduct a systematic investigation of training and inference scaling laws in generative retrieval, exploring how model size, training data scale, and inference-time compute jointly influence retrieval performance. To address the lack of suitable metrics, we propose a novel evaluation measure inspired by contrastive entropy and generation loss, providing a continuous performance signal that enables robust comparisons across diverse generative retrieval methods. Our experiments show that n-gram-based methods demonstrate strong alignment with both training and inference scaling laws, especially when paired with larger LLMs. Furthermore, increasing inference computation yields substantial performance gains, revealing that generative retrieval can significantly benefit from higher compute budgets at inference. Across these settings, LLaMA models consistently outperform T5 models, suggesting a particular advantage for larger decoder-only models in generative retrieval. Taken together, our findings underscore that model sizes, data availability, and inference computation interact to unlock the full potential of generative retrieval, offering new insights for designing and optimizing future systems. We release code at SLGR GitHub repository.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and ranking
@@ 1. Introduction
Document retrieval is a fundamental area in information retrieval, focusing on retrieving relevant documents from large-scale corpora in response to user queries. Early retrieval systems were built on term-based heuristic methods, such as TF-IDF (Salton et al., 1975) and BM25 (Robertson and Zaragoza, 2009), which rely on query and document term overlap. With the development of pre-trained language models, like BERT (Devlin et al., 2019), retrieval evolved into the dense retrieval paradigm, where queries and documents are mapped into a shared high-dimensional vector space, achieving advanced performance in document retrieval. Recently, with the rise of generative large language models (LLMs) (OpenAI et al., 2024; Touvron et al., 2023; Raffel et al., 2020), a new paradigm called generative retrieval has emerged. Instead of matching queries with documents, generative retrieval directly generates documents based on a given query. By reformulating the retrieval task as an autoregressive generation problem, generative retrieval indeed introduces a novel solution to the research field.
Figure 1. Scaling laws of n-gram-based methods in training and inference. (1) Training Scaling Law: Contrastive Generation Loss shows a power-law relationship with model size for both T5 and LLaMA models. (2) Inference Scaling Law: Miss Rate decreases consistently with increasing inference FLOPs, demonstrating the scalability of n-gram-based methods across the T5-Base and LLaMA-7B model.
A central challenge in generative retrieval lies in designing effective document identifiers to represent documents, as generating entire long documents is impractical. The current identifiers can be divided into two broad categories based on how they carry semantics. 1) Natural identifiers retain inherent semantic information by leveraging components like titles (De Cao et al., 2021; Lee et al., 2022b; Chen et al., 2022; Li et al., 2023a) or n-gram fragments (Bevilacqua et al., 2022; Chen et al., 2023; Li et al., 2023b; Wang et al., 2023) extracted from the original text. Titles provide concise, human-readable overviews, while n-gram snippets capture more granular semantic features. 2) Learned identifiers, on the other hand, derive semantic representations through clustering or codebook methods. Notable examples include numeric IDs (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2023a; Mehta et al., 2023) and codebook-derived tokens (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024; Zhang et al., 2024c), which discretize document embeddings into token sequences. Existing generative retrieval methods within these two categories have continued to evolve, showing promising performance.
@@
To address the above challenges, we introduce a new metric and employ larger models across different retrieval methods to systematically investigate how model size, training data scale, and inference compute impact performance. To capture subtle performance variations beyond discrete metrics, we propose a novel evaluation metric inspired by contrastive entropy and generation loss in neural scaling laws (Kaplan et al., 2020). Our metric measures the probability of generating the correct document identifier for a given query while considering random negative samples, yielding a continuous and sensitive performance signal. This metric enables consistent comparisons across various generative retrieval methods, models, and data scales. Leveraging this metric, we conduct extensive experiments to uncover the scaling behaviors of generative retrieval under different model sizes and data sizes. Additionally, we analyze how increased inference-time computation influences recall, providing further insight into the trade-offs and advantages of scaling computation.
From our extensive exploration, several intriguing findings stand out. 1) We observe that n-gram-based generative retrieval aligns remarkably well with both training and inference scaling laws, as illustrated in Figure 1, which presents clear scaling curves under varying model sizes and inference computation. 2) Expanding the training data scale benefits all methods, and n-gram-based approaches demonstrate especially robust gains, indicating a strong synergy between LLMs and natural identifiers. 3) We find that boosting inference computation (e.g., FLOPs) yields clear performance improvements that follow power-law trends, revealing that generative retrieval can significantly profit from additional inference computation—an aspect rarely discussed in prior work. 4) We discover that LLaMA models consistently outperform T5 models and exhibit higher theoretical upper bounds, hinting that the generative ability of larger decoder-only models may be particularly advantageous for generative retrieval.
@@ 2. Related Work
In this section, we revisit the background of generative retrieval and scaling laws. We first present the key advancements in generative retrieval, focusing on document identifier design and training methodologies. Then, we discuss neural scaling laws and their applications in various domains, including retrieval tasks.
@@ 2.1. Generative Retrieval
Generative retrieval leverages generative models, particularly LLMs, reformulating the retrieval task as an auto-regressive generation problem. Document identifiers, as the core component, should be distinctive enough to represent related passages, reducing useless information in a passage and making it easier for the model to memorize and learn. The design of DocID can be broadly categorized into two approaches: natural identifiers and learned identifiers.
The first type uses identifiers that naturally carry semantic meaning about its associated document, which are comparably cost-efficient to be created, without the need for additional human supervision or forcing any structure in the search space. A line of work (De Cao et al., 2021; Lee et al., 2022b; Chen et al., 2022; Li et al., 2023a) explored titles as identifiers and DSI(Tay et al., 2022) tested the first N words of the passage, providing human-readable summaries of document content. Multiple textual fragments  (Bevilacqua et al., 2022; Chen et al., 2023; Li et al., 2023b; Wang et al., 2023) extracted directly from the document are then proved to capture richer semantic features of the document. These n-grams are not tied to a specific document but reflect shared semantic content across multiple semantically related documents. The second type called learned identifiers, acquires semantic meaning through dense representation of passages clustering or codebook training, which induces structure in the search space with semantically similar documents having more similar document IDs. In this case, each document is assigned a unique numerical representation, and retrieval is achieved by generating the corresponding identifier. Specifically, numeric ID (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2023a; Mehta et al., 2023) was found easy to construct but required extra memory steps. While codebook-derived tokens (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024; Zhang et al., 2024c) have been proven effective because the search space is reduced after each decoding step.
Generative retrieval research has also advanced in training methods, which can be divided into generative training and discriminative training. In generative training, the model is trained to generate the appropriate identifier for a given query (Pradeep et al., 2023; Tay et al., 2022; Wang et al., 2022; Zeng et al., 2024; Bevilacqua et al., 2022), aligning naturally with the generative capabilities of LLMs to produce accurate document identifiers. Discriminative training, on the other hand, employs ranking losses (Li et al., 2024; Tang et al., 2024) and negative sample mining techniques (Zeng et al., 2024) to teach the model to produce a ranked list of documents. These approaches directly optimize retrieval performance by aligning the training objectives with the ranking requirements of retrieval tasks.
@@ 2.2. Neural Scaling Laws
Neural scaling laws describe predictable patterns of performance improvement as model size, dataset size, and computational resources increase. Baidu (Hestness et al., 2017) first introduced power-law relationships between test loss and these factors, offering an insight to predict neural network training. OpenAI (Kaplan et al., 2020) extended this concept to larger models, demonstrating that model scaling yields consistent improvements in tasks like language modeling. Google (Hoffmann et al., 2022) further advanced this understanding by introducing a unified formula for scaling laws, thus laying the groundwork for scaling strategies in neural networks.
Scaling laws have been successfully applied to various domain-specific fields such as speech recognition (Radford et al., 2023), computer vision (Dehghani et al., 2023; Zhai et al., 2022), and vision-language models (Jia et al., 2021; Radford et al., 2021). In the field of Information Retrieval, scaling laws have been explored in recommendation. Studies have examined applications in Click-Through Rate (CTR) prediction (Ardalani et al., 2022) and sequential recommendation models (Zhang et al., 2024a) using unique item identifiers. Recent research has demonstrated the effectiveness of trillion-parameter sequential transducers for generative recommendations (Zhai et al., 2025) and the development of architectures like Wukong (Zhang et al., 2024b) has established scaling laws for large-scale recommendation systems by effectively capturing diverse, high-order interactions through scalable network layers.
However, research on scaling laws in retrieval tasks remains limited. Scaling laws for dense retrieval (Fang et al., 2024) have been investigated, focusing on embedding-based methods using BERT-like models (Devlin et al., 2019). Another study on industrial multi-stage advertisement retrieval systems (Wang et al., 2024) emphasizes task-specific optimizations but lacks exploration of general scaling laws applicable to generative retrieval. In contrast, an early exploration of generative retrieval examines performance across varying corpus and model sizes (Pradeep et al., 2023) but is limited to learned identifier methods, leaving natural identifier methods unaddressed. Additionally, it only adopts different sizes of T5 models and does not fit power-law scaling relationships. These gaps motivate our comprehensive study, which examines generative retrieval under varying model sizes, data scales, and DocID strategies, aiming to shed light on broader scaling behaviors and guide future system design.
Following these discussions on generative retrieval and neural scaling laws, in this work, we conduct a comprehensive investigation of scaling behaviors in generative retrieval. By including larger decoder-only models and analyzing representative approaches, we aim to uncover key principles of scaling and provide insights to advance the design and optimization of generative retrieval systems.
@@ 3. Methodology
In this section, we first formalize the generative retrieval task in Section 3.1. Next, we describe the representative approaches for natural identifier and learned identifier: MINDER (Li et al., 2023b) and RIPOR (Zeng et al., 2024), respectively, in Section 3.2. The backbones and training configurations used for generative retrieval are introduced in Section 3.3. Finally, to address the limitations of traditional ranking metrics in reflecting the scaling behaviors of generative retrieval, we propose contrastive generation loss in Section 3.4.
@@ 3.2. Representative Generative Retrieval Methods
As outlined above, the document identifier is a crucial component of generative retrieval. There are two primary categories: natural identifiers and learned identifiers. In this study, we select representative techniques from each category to evaluate their scaling behaviors. For natural identifiers, we use the n-gram-based method as n-grams effectively capture diverse semantic relationships between queries and documents. For learned identifiers, we adopt a codebook-based approach, as it leverages advanced neural methods to encode semantic information more compactly and effectively.
@@ 3.2.1. N-gram-based Generative Retrieval
In this method, an LLM is trained to generate query-specific n-grams, which act as identifiers for ranking documents. The training process begins by selecting n-grams from documents based on their overlap or semantic similarity with user queries. These n-grams serve as the basis for training the LLM to predict relevant identifiers when presented with a query. At inference, the LLM generates n-grams for a query, and these are used to score and rank documents based on a heuristic function, such as n-gram frequency or semantic similarity.
For our experiments, we adopt the method of MINDER (Li et al., 2023b), focusing on extracting identifiers from the body text of documents.
@@ 3.2.2. Codebook-Based Generative Retrieval
Codebooks originate from techniques designed to create discrete visual data representations (Lee et al., 2022a; van den Oord et al., 2017). Learned codebooks for documents represent them as sequences of unique codes that effectively capture the semantics of their associated content (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024; Zhang et al., 2024c).
The process begins by encoding documents into dense vector representations using an encoder network. These dense vectors are then discretized into tokens by mapping them to entries in the learned codebook. Finally, a decoder network reconstructs the original document from the codebook tokens to ensure the generated representations are accurate and compact. The resulting code sequences serve as unique identifiers for documents, establishing a one-to-one correspondence between the sequence and the document. At inference, the LLM generates a code sequence for a query, which is matched to the corresponding document based on the learned codebook.
In our study, we select the RIPOR (Zeng et al., 2024) as the representative of the codebook-based method.
@@ 3.3.1. Backbone Models
For our experiments, we use the widely adopted T5 (Raffel et al., 2020) series as the primary backbone, which has been extensively employed in previous generative retrieval studies (Tay et al., 2022; Zeng et al., 2024; Ren et al., 2023; Mehta et al., 2023; Zhuang et al., 2023b; Zhang et al., 2024c). To evaluate the effect of model size, we experiment with all T5 variants: T5-Small, T5-Base, T5-Large, T5-XL, and T5-XXL, which differ only in parameter sizes while maintaining identical pre-training configurations.

### v2

@@ Abstract.
Generative retrieval reformulates retrieval as an autoregressive generation task, where large language models (LLMs) generate target documents directly from a query. As a novel paradigm, the mechanisms that underpin its performance and scalability remain largely unexplored. We systematically investigate training and inference scaling laws in generative retrieval, exploring how model size, training data scale, and inference-time compute jointly influence performance. We propose a novel evaluation metric inspired by contrastive entropy and generation loss, providing a continuous performance signal that enables robust comparisons across diverse generative retrieval methods. Our experiments show that n-gram-based methods align strongly with training and inference scaling laws. We find that increasing model size, training data scale, and inference-time compute all contribute to improved performance, highlighting the complementary roles of these factors in enhancing generative retrieval. Across these settings, LLaMA models consistently outperform T5 models, suggesting a particular advantage for larger decoder-only models in generative retrieval. Our findings underscore that model sizes, data availability, and inference computation interact to unlock the full potential of generative retrieval, offering new insights for designing and optimizing future systems. We release code at SLGR GitHub repository.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3729973isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Retrieval models and ranking
@@ 1. Introduction
Document retrieval is a fundamental area in information retrieval, focusing on retrieving relevant documents from large-scale corpora in response to user queries. Early retrieval systems were built on term-based heuristic methods, such as TF-IDF (Salton et al., 1975) and BM25 (Robertson and Zaragoza, 2009), which rely on query and document term overlap. With the development of pre-trained language models, like BERT (Devlin et al., 2019), retrieval evolved into the dense retrieval paradigm, where queries and documents are mapped into a shared high-dimensional vector space, achieving advanced performance in document retrieval. Recently, with the rise of generative large language models (LLMs) (OpenAI et al., 2024; Touvron et al., 2023; Raffel et al., 2020), a new paradigm called generative retrieval has emerged (Li et al., 2024a). Instead of matching queries with documents, generative retrieval directly generates documents based on a given query. By reformulating the retrieval task as an autoregressive generation problem, generative retrieval indeed introduces a novel solution to the research field.
Figure 1. Scaling laws of n-gram-based methods in training and inference. (1) Training Scaling Law: Contrastive Generation Loss shows a power-law relationship with model size for both T5 and LLaMA models. (2) Inference Scaling Law: Miss Rate decreases consistently with increasing inference FLOPs across the T5-Base and LLaMA-7B model.
A central challenge in generative retrieval lies in designing effective document identifiers to represent documents, as generating entire long documents is impractical. The current identifiers can be divided into two broad categories based on how they carry semantics. 1) Natural identifiers retain inherent semantic information by leveraging components like titles (De Cao et al., 2021; Lee et al., 2022b; Chen et al., 2022; Li et al., 2023a) or n-gram fragments (Bevilacqua et al., 2022; Chen et al., 2023; Li et al., 2023b; Wang et al., 2023) extracted from the original text. Titles provide concise, human-readable overviews, while n-gram snippets capture more granular semantic features. 2) Learned identifiers, on the other hand, derive semantic representations through clustering or codebook methods. Notable examples include numeric IDs (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2023a; Mehta et al., 2023) and codebook-derived tokens (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024b; Zhang et al., 2024c), which discretize document embeddings into token sequences. Existing generative retrieval methods within these two categories have continued to evolve, showing promising performance.
@@
To address the above challenges, we introduce a new metric and employ larger models across different retrieval methods to systematically investigate how model size, training data scale, and inference compute impact performance. To capture subtle performance variations beyond discrete metrics, we propose a novel evaluation metric inspired by contrastive entropy and generation loss in neural scaling laws (Kaplan et al., 2020). Our metric measures the probability of generating the correct document identifier for a given query while considering random negative samples, yielding a continuous and sensitive performance signal. This metric enables consistent comparisons across various generative retrieval methods, models, and data scales. Leveraging this metric, we conduct extensive experiments to uncover the scaling behaviors of generative retrieval under different model sizes and data sizes. Additionally, we analyze how increased inference-time computation influences performance, highlighting its role in improving retrieval accuracy.
From our extensive exploration, several intriguing findings stand out. 1) We observe that n-gram-based generative retrieval aligns remarkably well with both training and inference scaling laws, as illustrated in Figure 1, which presents clear scaling curves under varying model sizes and inference computation. 2) Expanding the training data scale benefits all methods, and n-gram-based approaches demonstrate especially robust gains, indicating a strong synergy between LLMs and natural identifiers. 3) We discover that LLaMA models consistently outperform T5 models and exhibit higher theoretical upper bounds, hinting that the generative ability of larger decoder-only models may be particularly advantageous for generative retrieval. 4) We find that boosting inference computation yields clear performance improvements that follow power-law trends, revealing that generative retrieval can significantly profit from additional inference computation—an aspect rarely discussed in prior work.
@@ 2. Related Work
In this section, we revisit the previous studies of generative retrieval and scaling laws. We first present the key advancements in generative retrieval, focusing on document identifier design and training methodologies. Then, we discuss neural scaling laws and their applications in various domains, including retrieval tasks.
@@ 2.1. Generative Retrieval
Generative retrieval reformulates retrieval as an autoregressive generation task, where a language model directly generates the identifier of the target document given a query. A central component in generative retrieval is the document identifier (DocID), which serves as the generation target. DocID can be broadly categorized into two types: natural identifiers and learned identifiers.
The first type uses identifiers that naturally carry semantic meaning about its associated document, which are comparably cost-efficient to be created, without the need for additional human supervision or forcing any structure in the search space. A line of work (De Cao et al., 2021; Lee et al., 2022b; Chen et al., 2022; Li et al., 2023a) explored titles as identifiers and DSI (Tay et al., 2022) tested the first N words of the passage, providing human-readable summaries of document content. Multiple textual fragments  (Bevilacqua et al., 2022; Chen et al., 2023; Li et al., 2023b; Wang et al., 2023) extracted directly from the document are then proved to capture richer semantic features of the document. These n-grams are not tied to a specific document but reflect shared semantic content across multiple semantically related documents. The second type called learned identifiers, acquires semantic meaning through dense representation of passages clustering or codebook training, which induces structure in the search space with semantically similar documents having more similar document IDs. In this case, each document is assigned a unique numerical representation, and retrieval is achieved by generating the corresponding identifier. Specifically, numeric ID (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2023a; Mehta et al., 2023) was found easy to construct but required extra memory steps. And codebook-derived tokens (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024b; Zhang et al., 2024c) have been proven effective because the search space is reduced after each decoding step.
Generative retrieval research has also advanced in training methods, which can be divided into generative training and discriminative training. In generative training, the model is trained to generate the appropriate identifier for a given query (Pradeep et al., 2023; Tay et al., 2022; Wang et al., 2022; Zeng et al., 2024b; Bevilacqua et al., 2022), aligning naturally with the generative capabilities of LLMs to produce accurate document identifiers. Discriminative training, on the other hand, employs ranking losses (Li et al., 2024b; Tang et al., 2024) and negative sample mining techniques (Zeng et al., 2024b) to teach the model to produce a ranked list of documents. These approaches align the training objectives with the ranking requirements of retrieval tasks.
@@ 2.2. Neural Scaling Laws
Neural scaling laws describe predictable patterns of performance improvement as model size, dataset size, and computational resources increase. Baidu (Hestness et al., 2017) first introduced power-law relationships between test loss and these factors, offering an insight to predict neural network training. OpenAI (Kaplan et al., 2020) extended this concept to larger models, demonstrating that model scaling yields consistent improvements in tasks like language modeling. Google (Hoffmann et al., 2022) further introduced a unified formula for scaling laws, thus laying the groundwork for scaling strategies in neural networks.
Scaling laws have been successfully applied to various domain-specific fields such as speech recognition (Radford et al., 2023), computer vision (Dehghani et al., 2023; Zhai et al., 2022), and vision-language models (Jia et al., 2021; Radford et al., 2021). In the field of information retrieval, scaling laws have been explored in recommendation (Lin et al., 2024). Studies have examined applications in Click-Through Rate (CTR) prediction (Ardalani et al., 2022) and sequential recommendation models (Zhang et al., 2024a) using unique item identifiers. Recent research has demonstrated the effectiveness of trillion-parameter sequential transducers for generative recommendations (Zhai et al., 2025) and the development of architectures like Wukong (Zhang et al., 2024b) has established scaling laws for large-scale recommendation systems by effectively capturing diverse, high-order interactions through scalable network layers.
However, research on scaling laws in retrieval tasks remains limited. Scaling laws for dense retrieval (Fang et al., 2024) have been investigated, focusing on embedding-based methods using BERT-like models (Devlin et al., 2019). Another study on industrial multi-stage advertisement retrieval systems (Wang et al., 2024) emphasizes task-specific optimizations but lacks exploration of general scaling laws applicable to generative retrieval. An early exploration of generative retrieval examines performance across varying corpus and model sizes (Pradeep et al., 2023) but is limited to learned identifier methods, leaving natural identifier methods unaddressed. Additionally, it only adopts different sizes of T5 models and does not fit power-law scaling relationships.
In contrast to prior work, our study investigates generative retrieval with larger decoder-only models, compares both learned and natural DocID strategies, and examines whether power-law relationships hold across model and data scales. Moreover, we analyze inference-time scaling behavior, which has received little attention in previous studies. These distinctions enable a more comprehensive understanding of generative retrieval scaling beyond existing task-specific or partial analyses.
@@ 3. Methodology
In this section, we first formalize the generative retrieval task in Section 3.1. Next, we describe the representative approaches for natural identifier and learned identifier: MINDER (Li et al., 2023b) and RIPOR (Zeng et al., 2024b), respectively, in Section 3.2. The backbones and training configurations used for generative retrieval are introduced in Section 3.3. Finally, to address the limitations of traditional ranking metrics in reflecting the scaling behaviors of generative retrieval, we propose contrastive generation loss in Section 3.4.
@@ 3.2. Representative Generative Retrieval Methods
As outlined above, the document identifier is a crucial component of generative retrieval. There are two primary categories: natural identifiers and learned identifiers. In this study, we select representative techniques from each category to evaluate their scaling behaviors. For natural identifiers, we use the n-gram-based method as n-grams effectively capture diverse semantic relationships between queries and documents. For learned identifiers, we adopt a codebook-based approach, as it leverages advanced neural methods to encode semantic information effectively.
@@ 3.2.1. N-gram-based Generative Retrieval
In this method, an LLM is trained to generate query-specific n-grams, which act as identifiers for ranking documents. As shown in Figure 2 (1), the training process begins by selecting n-grams from documents based on their overlap or semantic similarity with user queries. These n-grams serve as the basis for training the LLM to predict relevant identifiers given a query. At inference, the LLM generates n-grams for a query, and these are used to score and rank documents based on a heuristic function, such as n-gram frequency or semantic similarity. For our experiments, we adopt the method of MINDER (Li et al., 2023b), focusing on extracting identifiers from the body text of documents.
@@ 3.2.2. Codebook-Based Generative Retrieval
Codebooks originate from techniques designed to create discrete visual data representations (Lee et al., 2022a; van den Oord et al., 2017). Learned codebooks for documents represent them as sequences of unique codes that effectively capture the semantics of their associated content (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024b; Zhang et al., 2024c).
As shown in Figure 2 (2), the process begins by encoding documents into dense vector representations using an encoder network. These dense vectors are then discretized into tokens by mapping them to entries in the learned codebook. Finally, a decoder network reconstructs the original document from the codebook to ensure the generated representations are accurate and compact. The resulting code sequences serve as unique identifiers for documents, establishing a one-to-one correspondence between the code sequence and the document. At inference, the LLM generates a code sequence for a query, which is matched to the corresponding document based on the learned codebook. In our study, we select the RIPOR (Zeng et al., 2024b) as the representative of the codebook-based method.
Figure 2. Comparison of n-gram-based and codebook-based methods. N-grams are extracted text spans, while codebook-based methods are discrete representations generated via codebook.
@@ 3.3.1. Backbone Models
For our experiments, we use the widely adopted T5 (Raffel et al., 2020) series as the primary backbone, which has been extensively employed in previous generative retrieval studies (Tay et al., 2022; Zeng et al., 2024b; Ren et al., 2023; Mehta et al., 2023; Zhuang et al., 2023b; Zhang et al., 2024c). To evaluate the effect of model size, we experiment with all T5 variants: T5-Small, T5-Base, T5-Large, T5-XL, and T5-XXL, which differ only in parameter sizes while maintaining identical pre-training configurations.